Presentation is loading. Please wait.

Presentation is loading. Please wait.

1 Update on project management for MICE Project Status and schedule Richard Apsimon MICE CMPB 5 th Nov 2007.

Similar presentations

Presentation on theme: "1 Update on project management for MICE Project Status and schedule Richard Apsimon MICE CMPB 5 th Nov 2007."— Presentation transcript:

1 1 Update on project management for MICE Project Status and schedule Richard Apsimon MICE CMPB 5 th Nov 2007

2 2 Project Management The departure of Paul Drumm left a very sizable hole which Ken has filled for several months. More recently, Andy Nichols (STBU) and Richard Apsimon (PPD) have joined to take a joint Technical Coordinator/ Project Manager role. Willie Spensley is MICE Hall manager and Principal Contractor. Ashok Jamdagni (IC) and Paul Flower (PPD) and the Hall Foremen. Chris Rogers is the ISIS Liaison officer Malcolm Ellis (Brunel) is the MICE Integration Physicist …and we have volunteers to fill the roll of MICE Operation Manager (MOM) i.e. Run Coordinator.


4 4 Weekly MICE Hall planning meetings. Essentially a 2 week look-ahead Weekly MICE Schedule planning meetings, concentrating On the longer term plus planning for the forthcoming weekend

5 5 MICE Target Previous problems with bearing wear. Now using DLC (Diamond Like Carbon or Diamolith). It is a Plasma Enhanced Chemical Vapour Deposition (PECVD) process that deposits a material layer that is about 3-5um thick. At present only got one, top bearing, DLC coated and the target shaft. >3M operations without wear

6 6 MICE Target Problems that we needed to solve for target installation into ISIS: –Removal of old target….. NOW during Xmas shutdown Hoped to demonstrate procedure for removal of hot target BUT we wanted to use short Nov. shutdown ISIS RPA has agreed that we dont need to go through the full procedure. We planned removal this week NO LONGER POSSIBLE – need 1 week cool-down –200µSv/h in working area, –2500µSv/h on the beam pipe, –~400µSv/h near the target bellows/beam pipe interface –500µSv/h on the downstream bellows –Approval to put the new target into ISIS. YES! MICE procedure accepted by ISIS –Some conditions –Infrastructure ready for target in ISIS. Two targets needed, test vessel, remote handling, fibre optics, control cables. Bottom bearings need coating

7 7 Target: schedule:

8 8 Target: schedule continued: Now planned with DL controls group: needs to be updated.

9 9 Upstream beamline Require to complete/install: –Alignment PROBLEM – magnet stand SOLUTION – New support stand to be built in time for Xmas shutdown –Vacuum vessels PROBLEM - with Dipole 1 Vacuum box in wrong place. PROBLEM - with ISIS beam-pipe beam would hit 25mm of steel. SOLUTION – Manufacture paused while box design altered to move thin window. –Services: Vacuum, Water & Electrical –Agreement with M.Krendler (ISIS): –Milestone (deadline) for completion of above - 20 Jan08 Working with individuals to complete as much as possible before by using the current minor maintenance period

10 10 Upstream beam line: power supplies: Rack being installed now: –Required removal of treads of north-east stair case Power-supply units and cabling to follow

11 11 Upstream beam line: Xmas work

12 12 DSA (Decay Solenoid Area) June: Hole to synchrotron vault, North wall in place October: Shielding to synchrotron, W & S walls of iron & concrete built. PPS doors installed. Plinth built, additional shielding on plinth, wall painted. Dipole 2 installed, First Q35 baseplate installed, surveyed & grouted.

13 13 View over DSA into MICE Hall June: Most of Hall cleared, some shielding October: Decay Solenoid Area civil engineering almost complete. Roof still to be designed & fitted. This week, we plan to remove sandbags from synchrotron hall side and replace with bags of steel shot. This should allow us to remove North wall before Xmas.

14 14 Delay Solenoid Major leak has been found and repaired BUT at the cost of considerable damage to MLI. Need 2? weeks of skilled effort. Matt Hills is developing a plan to complete this task. Cold mass still to be aligned and tie rods re- tensioned. We are working hard to find the correct technical help for this work.

15 15 Linde cryogenics After DSA completed, support installed for buffer tank, below which is control plant. Transfer line carries cold helium gas to and from either Test Cap or decay solenoid

16 16 Linde cryogenics Helium transfer line installed and connected to Test Cap. Most of the scaffolding now removed. After cool-down and other tests, transfer line will be connected to PSI Decay Solenoid Installation completed ahead of schedule Commissioning started and hope to finish 25 th or even 18 th Nov.

17 17 Linde cryogenics - Update Latest news: –Heat leak located in flexible section of transfer line. Insufficient cooling available at the test cap. DeMaCo engineers arrived on Saturday to dismantle & inspect the transfer line Discovered that the flexible section had been constructed to the wrong dimensions They suggested a temporary fix on site, to allow commissioning with the test-cap to continue –This was rebuffed by Christian Dachauer of Linde and by Mike Courthold Therefore, the transfer line was taken back to Holland on Sunday morning by the DeMaCo engineers for re-engineering It will be brought back to RAL and refitted on Thursday 8/11/07

18 18 Linde Schedule as of 9 th Oct

19 19 Q35 Refurbishment & Installation First 3 magnets finished this week. Test in R6. One technician has left. Replacement in place. Need to increase # people with crane licences –Trial assembly of magnet support on baseplate complete –Upstream baseplate (Q4-Q6) in DSA has been aligned. Levelled & grouted. Downstream plate ready with all components and will be installed as soon as Linde equipment & spares can be moved.

20 20 Down-stream beam-line: Quadrupole triplets: –First quad triplet refurbished: Magnetic testing to start next week –Trial build of quad stand complete Quadrupole triplets cont.: –Refurbishment of remaining quads: Goal: complete by 20Jan08 –Services being installed now

21 21 Q35 refurbishment: schedule: Focus on Q35 work has paid off, but: –Magnetic testing will start 01Nov07: Developing schedule with Paul Flower now he has returned from CERN –Have had one more coil failure: potential issue: Have expedited testing of remaining coils Continuing refurbishment –Need to consider options when it is known if any more coils have failed

22 22 MICE Local Control Room (MLCR) Rack bases delivered DPM & floor screed completed Network & PPS racks, populated and tested (DL) and delivered last week. Remaining racks last/this week. Air conditioning and power to start soon. –Estimated to take three weeks so complete by 9th November. Roof has been fitted and plastered and rack bases and racks will move in from mid-week. Network connection cabling installed. –Connection to RAL site to star soon –A wireless access point to be set up as soon as possible. Further wiring into the hall to follow. –Still needs to be coordinated with the civil engineering that is going on.

23 23 MLCR timescale

24 24 Reviews: beam line

25 25 Spectrometer and instrumentation: Spectrometer solenoid: –First spectrometer solenoid: Coils almost complete Expect delivery to FNAL for magnetic field mapping in December –Second spectrometer solenoid: Fabrication advanced Expect delivery ~2 months after first CKov: –Two modules: Most of the first (CKova) at RAL Second module under construction in Mississippi TOF: –TOF0 and TOF1 under construction in Milan: Due at RAL 15 th November for cosmic ray tests BUT not expected until around 15 th Dec. Calorimeter: –Pre-sampler (KL): under construction in Rome Also due at RAL 15 th November for cosmic ray tests –Sandwich calorimeter: Design complete, prototype under construction

26 26 First full tracker being manufactured: –Final step in QA chain (station acceptance rig) now in full production mode –Space frame being built –Support frame in manufacture Electronics: –Being prepared for shipping in FNAL –Will be shipped this week Spectrometer and instrumentation: continued

27 27 Cosmic-ray integration test: Will start 12Nov07: –Will include: CKov, Tracker, ToF, and KL (pre- sampler) 18/10/07 Proposed lower Cosmictest stand: plan (using approximate dimensions) D2 CKOV KL (1.5 x2.1M) 90mmSq? crossbeams support components

28 28 Cosmic-ray integration test: The cosmic ray test was initially planned to be held in R12, but this proved to be difficult due to space constraints. We have now been given some space in R8 which we have until the end of January. Work at RAL currently to sort out: (Unforeseen expense (likely small, but being estimated) Mains and 3 phase power Water supply for compressors Network Safety interlock systems for the tracker readout A tent has been identified that is the right size and has material which has been tested to exclude the dangerous UV light.

29 29 Other Issues Xmas shutdown –Well planned, WBS manager (Tim Hayler) –Already VERY tight. Every other task that is delayed now moves into this schedule. (e.g Target removal, upstream beamline, ISIS vacuum vessels etc.) False floor and plinth in main hall –It might be possible to complete false floor in time for decay solenoid to move for cool-down tests -UNLIKELY –Otherwise we must build a temporary support Other news –Water chiller order placed. –Rolling platform FDR completed, PRR due. –Magnetic shield wall drgs. One wall out for tender. Aim to have contracts placed well before end of this FY. –Quotes for Cyclotron steel being sought –Cost for both items likely to be higher than anticipated.

30 30 Phase I instrumentation: –Components: Tracker: –Station acceptance QA rig now running at Imperial »Station 5 (first final station) has been QAed: results OK; »Decided last week(?): proceed with QA or more studies required on station 5 –Space frame in manufacture –Cosmic stand signed off for manufacture (?) TOF & KL: –Due to arrive 15Nov07 - LATE CKova –Already at RAL, needs to be assembled DAQ –Already at RAL, needs to be moved to R8 –Should be supervised by DAQ-expert

31 31 Cooling channel components: RFCC module: –CC design and fabrication done in collaboration by LBNL and ICST in Harbin, China conductor ready to order –RF cavities will be similar to existing MuCool prototype fabrication to get under way shortly –MuCool 201 MHz cavity R&D: Without field now have 19 MV/M

32 32 Cooling channel components: AFC module: Focus coil module presently out for tender –anticipate contract award in September 2007 –two coils that can run with same or opposite polarity »20-L LH 2 absorber (plus safety windows) fits inside Oxford, RAL, KEK

33 33 Current concerns Preparation of –The PSI solenoid – MLI insulation –MICE target –Upstream beamline Is the misalignment critical? YES Can it be solved in time? YES –Connection of services for upstream & downstream beamline components –Transfer line –Remaining Q35s. Can we finish by 20/1/08? Radiation surveys –Many areas where these matter; Future target removal Connection of magnets in synchrotron vault - OK Work inside DSA – removal of North wall –We hope the steel shot will solve this one Excavation of earth for H2 vents and pumps

Download ppt "1 Update on project management for MICE Project Status and schedule Richard Apsimon MICE CMPB 5 th Nov 2007."

Similar presentations

Ads by Google